One number is 9 times a first number. A third is 100 more than the first number. If the sum of the three numbers is 518, find the numbers.
Answer by macston(5194)

.
F=first number; S=second number=9F; T=third number=F+100
.
F+S+T=518
F+9F+F+100=518
11F=418
F=38
S=9F=342
T=F+100=38+100=138
.
ANSWER: The numbers are 38, 342, and 138.
.
CHECK:
F+S+T=518
38+342+138=518
518=518
.